OpenStudy (anonymous):

What is the area in square millimeters of a rectangle with a width of 8.7 centimeters, and a length of 9.3 centimeters? A. 8091 mm2 B. 809.1 mm2 C. 80,910 mm2 D. 80.91 mm2

OpenStudy (sleepyhead314):

Hello @KatnissMallark :) you are currently in the Biology section you would get a better and faster answer if you would ask this in the Mathematics section :) the Area of a Rectangle = (length) * (width) but they want the answer in "millimeters" although they gave you the length and stuff in "centimeters" there are 10 millimeters in 1 centimeter so to convert centimeter to millimeters you would multiply the given numbers by 10 so width = 8.7cm = 8.7*10 = 87 mm length = 9.3cm = 9.3 * 10 = 93 mm area of a rectangle = (87mm) * (93mm) = ?